# This Candidate is [MASK] Prompt-based Sentiment Extraction and Reference Letters

### v2

@@ Abstract
I propose a relatively simple way to deploy pre-trained large language models (LLMs) in order to extract sentiment and other useful features from text data. The method, which I refer to as prompt-based sentiment extraction, offers multiple advantages over other methods used in economics and finance. I apply my prompt-based strategy to a hand-collected corpus of confidential reference letters (RLs). I show that the sentiment contents of RLs is clearly reflected in job market outcomes. Candidates with higher average sentiment in their letters perform markedly better regardless of the measure of success chosen. Moreover, I show that disagreement among letter writers negatively affects the job market candidate’s performance. I compare my sentiment extraction approach to other commonly used methods for sentiment analysis: ‘bag-of-words’ approaches, fine-tuned language models, and querying advanced chatbots. I find that no other method can reproduce the results obtained by prompt-based sentiment extraction. Finally, I slightly modify the method to obtain ‘gendered’ sentiment scores (as in Eberhardt et al., 2023). I show that letters of reference written for female candidates emphasize ‘grindstone’ personality traits, whereas male candidates’ letters emphasize ‘standout’ traits. These gender differences negatively affect women’s job market outcomes.
@@ 1 Introduction
I apply prompt-based sentiment extraction to a hand-collected corpus of confidential RLs submitted as a standard part of the package that academic economists submit to potential employers in academia and other sectors. Letters of recommendation have a number of desirable features that make them an ideal setting to test my sentiment extraction method. Because they are confidential, we can be sure that they have not been used as part of the training set for any of the open-source LLMs that I test-drive. As argued recently by Ludwig et al. (2025), the existence of leakage between the LLM’s training data and the economic application has the potential to severely bias results. In addition, RLs provide an interesting mix of relatively objective/factual descriptions and more opinionated content. While the structure of the letters is relatively stable, the opinions and discussion can often be quite idiosyncratic, so the task of producing a sentiment score is far from trivial. Finally, on average letters are around 2,000 words long, which is relatively easy to process for state-of-the-art LLMs.
@@
There is also some experimental evidence on the value of reference letters. Pallais (2014) conducted a field experiment on an online marketplace in which workers bid to be hired for short term employment. One of her treatment groups received detailed feedback after completing a data-entry task, while another group received only a one-to-five rating. She found that, in the two months after treatment, both treatment groups benefitted from treatment relative to a control group that was not hired. Abel et al. (2020) conducted an audit study where they randomly appended a reference letter to some applications submitted on behalf of job seekers to vacancies in South Africa. They found that, for the same applicant, attaching a letter increased the probability of receiving a response and the rate of interview requests by about 60 percent.
@@
In summary, I find that prompt-based sentiment extraction produces sentiment scores that are strongly correlated with the probability of obtaining a good job market outcome. In particular, high average sentiment among all referees are predictive of a sizeable higher probability of success. Moreover, my paper is the first to document that disagreement among letter writers can offset the positive effect of high average sentiment, which is strictly consistent with employers taking all available information in RLs pretty much at face value. While these findings are not causal, they strongly suggest that the theoretical literature might have over-emphasized the potential pitfalls of reference letters.
@@ 2.1 Lexical Approach
A final limitation of the lexical approach is that, if the method is to be used to generate a measure other than a sentiment score, a completely new dictionary must be produced. For example, Schmader et al. (2007) created a set of dictionaries to summarize different semantic aspects of reference letters written for chemists and biochemists. Specifically, they developed sets of terms associated with the following language categories: grindstone traits, ability traits, research terms, teaching and citizenship terms, and standout adjectives. Eberhardt et al. (2023) applied a similar strategy to letters of recommendation written on behalf of job market candidates applying to a top UK economics department. [^10] A non-trivial additional layer of complexity involves how to normalize the term counts in non-sentiment oriented dictionary tasks. Schmader et al. (2007) consider the fraction of words in a letter that belong to each of their dictionaries. Eberhardt et al. (2023) use the term-frequency-inverse-document-frequency (tf-idf) when computing the scores associated with each language category.
@@ 2.3.2 Model Choice
All the LLMs mentioned so far belong to the foundational class, meaning that they have not been fine-tuned to follow human instructions or to produce output that is “aligned” with the objectives of humans. Instruct models, like Llama-instruct (Touvron et al., 2023), GPT-instruct (Ouyang et al., 2022), or DeepSeek-V3 (DeepSeek-AI et al., 2025) are “post”-trained using supervised data of question-answer pairs, as well as other techniques like reinforcement learning with human feedback, in which the model produces several alternative outputs that a human user later ranks. [^24] World famous applications like ChatGPT add a memory layer (and other features) to the otherwise stateless instruct model.
@@
It is important to reiterate that the “score” that results from these instructions is quite unlike the polarity score from equation (2). The former consists of the sequence of text tokens making up a floating point number within -1 and 1 that results from the instruct model’s sampling process. The latter is a summary measure of the whole probability distribution over the foundational model’s vocabulary.
@@ 3.2 Reference Letter Sentiment
Figure 1 (left) shows the kernel estimate of the distribution of prompt-based polarity scores for the full sample of RLs. Several facts are apparent. Sentiment polarity is overwhelmingly positive; there are only a couple of letters with negative polarity. The distribution of sentiment is roughly symmetric, with mean sentiment of about 0.07.303030Table A.1 in the appendix has descriptive statistics for all the sentiment measures used in this paper.
@@
Letter sentiment tends to reinforce other known signals of higher ability. Candidates from PhD programs in America and from programs in the top 25 have substantially better RLs. The same is true about candidates with top publications, as well as candidates who managed to procure a letter from a top-5 percentile research economist. These statistics are in line with what a general sentiment measure should look like under the general assumption that higher ability candidates get more enthusiastic references. In section 5 below, I further analyze the determinants of sentiment polarity as well as other RL features extracted via prompts.
@@ 4 Predicting Job Market Outcomes
There are multiple potentially appropriate measures of what constitutes “success” in the academic job market. Job market outcomes are heterogenous on multiple fronts. Consider figure 3, depicting different aspects of candidates’ placement by September after the job market season. Outcomes vary by the sector of employment and the type of position. In addition, employers differ on how they rank according to their research output, with higher ranked employers being markedly more desirable to the average junior academic. [^31] In order to determine the employer’s ranking, I once again use Tilburg’s ranking for academic institutions. For policy institutions, I use the econphd.net ranking (https://econphd.econwiki.com/rank/rallec.htm), which focuses on publications and has a transparent methodology.
@@
In what follows, I define my baseline “success” outcome as attaining a top-100 academic or top-50 policy job. For some clarity on these cutoffs, consider that economics departments ranked 98–102 in the Tilburg ranking are NHH (Norwegian School of Economics), INSEAD, the California Institute of Technology and the University of Gothenburg. Institutions ranked in the top-50 of the policy ranking include the Federal Reserve Banks and the Board of Governors, as well as the ECB and a few other major European central banks, plus the World Bank, the IMF and a couple other major development institutes. In figure 3, the baseline cutoffs are depicted by vertical red lines. Clearly, only a minority of candidates in my sample attained success based on this measure.
@@
Baseline Top-100 Academic Employers & Top-50 Policy Institutions 175 27.13
Alt. 1 Exclude teaching positions (liberal arts colleges) and post-docs from the success group 115 17.83
Alt. 2 Add some highly paid tech jobs (Amazon, Meta, etc.) to the success group 190 29.46
Alt. 3 Use somewhat more forgiving cutoffs: top-150 academic and top-75 policy 243 37.67
Alt. 4 Use somewhat more restrictive cutoffs: top-75 academic and top-35 policy 150 23.26
@@
Notes: NN=645. $N_{success}$ shows by how many alternative definitions the candidate had success in the job market.
Figure 4 depicts the relationship between average RL quality and success in the academic job market. It is not hard to see that candidates who succeeded according to most definitions are clustered above and to the right relative to those who did not succeed.
@@ 4.1 Linear Regressions
$y_{it}=$ $\>\alpha+\beta_{1}\,\mathrm{avg\_RL\_length}_{it}+\beta_{2}\,\mathrm{avg\_RL\_sentiment}_{it}+\boldsymbol{X}^{\prime}_{it}\boldsymbol{\gamma}+\lambda_{t}+\epsilon_{it}$ (3)
where $y_{it}$ is the job market outcome of candidate $i$, in job market year $t$. The main predictors in these models are the average RL length (in thousands of words) and average RL sentiment polarity (in p.p.). [^32] In the interest of full clarity, the regressors are the average RL word count divided by 1,000 and the average sentiment polarity score multiplied by 100. These adjustments ease the interpretation of the coefficients by eliminating unnecessary decimal points. All regressions include controls for the number of RLs submitted on behalf of the candidate. Depending on the exact specification, the vector of control variables may also include: the candidate’s individual characteristics (sex, region of origin, field of research, publications, long PhD), the PhD program’s characteristics (rank, region), the job market year, and whether the letter writers are senior (specifically, the number of top-5 writers).
RLs are only one input in the recruitment process. Job market outcomes also depend on how the candidates present themselves in interviews, seminar presentations, and one-on-one meetings with faculty and administrators at potential employers, among other variables. Because these other determinants remain unobserved, the regression estimates cannot be interpreted causally. [^33] Assuming strong candidates get good letters and also tend to perform well in interviews and seminars, one reasonable expectation would be for regression coefficients to over-estimate the true causal effect that RL sentiment might have on a candidate’s probability of success.
However, in light of the theoretical discussion regarding whether referees can be trusted to offer honest assessments and whether hiring committees are properly incentivized to select the best candidates, it is still interesting to determine whether a statistically significant association exists between good letters and good job market outcomes. In this section and the next, I show that not only do higher average sentiment RLs are predictive of better outcomes, but also that the degree of disagreement among the letter writers—sentiment dispersion—is negatively associated with success in the market.
@@
Average Letter Length 0.2036∗∗∗ 0.2510∗∗∗ 0.2371∗∗∗ 0.2416∗∗∗ 0.2447∗∗∗
(000s words) (0.0473) (0.0469) (0.0470) (0.0503) (0.0506)
[0.0495] [0.0496] [0.0501] [0.0548] [0.0546]
||0.0470|| ||0.0450|| ||0.0446|| ||0.0501|| ||0.0515||
Average Letter Sentiment 0.0304∗∗∗ 0.0276∗∗∗ 0.0230∗∗∗ 0.0233∗∗∗ 0.0227∗∗∗
($p.$) (0.0071) (0.0067) (0.0069) (0.0070) (0.0070)
[0.0067] [0.0062] [0.0063] [0.0064] [0.0064]
@@
Table 7 shows regression estimates of equation (3) for the baseline outcome and with different sets of control variables. [^34] Table A.2 in the appendix presents the coefficient estimates for the control variables. Regardless of the specifics, both RL length and sentiment have a strong and highly statistically significant relationship with the probability of success. [^35] Throughout the paper I present standard errors, and the corresponding levels of significance, for three different assumptions regarding the joint distribution of the error terms in equation (3). HC0 standard errors assume the $\epsilon_{it}$ are independently distributed but heteroskedastic. Clustered standard errors allow for dependence across errors either within PhD programs (due to norms regarding how to write RLs, for example) or within clusters defined by the market year and the PhD rank group (due to competition among candidates for similar jobs). Interestingly, all the evidence suggests that the length of the RLs acts as an implicit positive signal in and of itself. A one-thousand word delta in average RL length is predictive of a very sizeable increase in the expected probability of success.
Prompt-based sentiment extraction enables a straightforward interpretation of the regression results, even if they cannot be taken to measure a causal effect. For example, an increase in average polarity of 0.1 means that the LLM assigns a 10 pp. higher probability to filling the prompt with one of the positive sentiment tokens relative to the negative tokens. According to the findings in the table, such as increase is predictive of a much higher ($\sim$23 p.p.) expected probability of obtaining a position in a top research job either in academia or in a policy institution.
@@ 4.2 Sentiment Dispersion
An open question is whether disagreement among the many letter writers is reflected in worse job market outcomes for the candidates. In order to study this issue, I focus on a sub-sample of “complete applications,” i.e. candidates for whom I have at least three RLs including the letter from the main research adviser. [^36] Table A.3 in the appendix replicates the analysis in table 8 for this sub-sample. The results are practically identical. Figure A.1 in the appendix shows that the distributions of sentiment polarity scores for different letter writers are almost identical. [^37] In unreported analysis, I found that the adviser’s RL sentiment and other writer’s RL sentiment have undistinguishable effects on job market outcomes. However, within-candidate variation in RL sentiment is quite high, accounting for over 45% of the total variance of sentiment. [^38] An alternative way to illustrate this point is by looking at correlations across letter writer scores. Excluding the fourth letter, which is available only for a minority of candidates, the correlations range between 0.27 and 0.40. In other words, letter writers tend to be aligned but are far from unanimous.
@@
Notes: NN=553 (complete applications sub-sample). The range is the difference between the max letter writer score and the min letter writer score for each candidate.
Table 9 shows summary statistics for alternative ways to measure sentiment dispersion among letter writers. The sentiment range—the measure that I will use as a baseline—is defined as the difference in sentiment scores between the writer with the highest score and the writer with the lowest score. For robustness, I will also consider the mean absolute deviation (MAD) and the standard deviation (SD) of scores across letter writers. As is clear from the table, all measures of dispersion are quite high relative to the average. They also vary significantly across candidates.
@@
$y_{it}=$ $\ \>\alpha+\beta_{1}\,\mathrm{avg\_RL\_length}_{it}+\beta_{2}\,\mathrm{avg\_RL\_sentiment}_{it}+\beta_{3}\,\mathrm{sentiment\_dispersion}_{it}$ (4)
$\>+\boldsymbol{X}^{\prime}_{it}\boldsymbol{\gamma}+\lambda_{t}+\epsilon_{it}$
where $\mathrm{sentiment\_dispersion}_{it}$ is one of the measures of disagreement across letter writers. Table 10 confirms that, regardless of the outcome definition, sentiment range has a negative relationship to success in the job market. Table A.4 in the appendix shows similar results for alternative measures of sentiment dispersion. In summary, these estimates suggest that RL sentiment has a determinant effect on job market outcomes and that hiring committees fully exploit the information contained in RLs.
@@
Four Letters & Avg. Letter Length Yes Yes Yes Yes Yes

### v3

@@ Abstract
I propose a relatively simple way to deploy pre-trained large language models (LLMs) in order to extract sentiment and other useful features from text data. The method, which I refer to as prompt-based sentiment extraction, offers multiple advantages over other methods used in economics and finance. In particular, it accepts the text input as is (without pre-processing) and produces a sentiment score that has a probability interpretation. Unlike other LLM-based approaches, it does not require any fine-tuning or labeled data. I apply my prompt-based strategy to a hand-collected corpus of confidential reference letters (RLs). I show that the sentiment contents of RLs are clearly reflected in job market outcomes. Candidates with higher average sentiment in their RLs perform markedly better regardless of the measure of success chosen. Moreover, I show that sentiment dispersion among letter writers negatively affects the job market candidate’s performance. I compare my sentiment extraction approach to other commonly used methods for sentiment analysis: ‘bag-of-words’ approaches, fine-tuned language models, and querying advanced chatbots. No other method can fully reproduce the results obtained by prompt-based sentiment extraction. Finally, I slightly modify the method to obtain ‘gendered’ sentiment scores (as in Eberhardt et al., 2023). I show that RLs written for female candidates emphasize ‘grindstone’ personality traits, whereas male candidates’ letters emphasize ‘standout’ traits. These gender differences negatively affect women’s job market outcomes.
@@ 1 Introduction
I apply prompt-based sentiment extraction to a hand-collected corpus of confidential RLs submitted as a standard part of the package that academic economists send to potential employers in academia and other sectors. Letters of recommendation have a number of desirable features that make them an ideal setting to test my sentiment extraction method. Because they are confidential, we can be sure that they have not been used as part of the training set for any of the open-source LLMs that I test-drive. As argued recently by Ludwig et al. (2025), the existence of leakage between the LLM’s training data and the economic application has the potential to severely bias results. In addition, RLs provide an interesting mix of relatively objective/factual descriptions and more opinionated content. While the structure of the letters is relatively stable, the opinions and discussion can often be quite idiosyncratic, so the task of producing a sentiment score is far from trivial. Finally, on average letters are around 2,000 words long, which is relatively easy to process for state-of-the-art LLMs.
@@
There is also some experimental evidence on the value of reference letters. Pallais (2014) conducted a field experiment on an online marketplace in which workers bid to be hired for short term employment. One of her treatment groups received detailed feedback after completing a data-entry task, while another group received only a one-to-five rating. She found that both treatment groups benefitted from treatment relative to a control group that was not hired. Abel et al. (2020) conducted an audit study where they randomly appended a reference letter to some applications submitted on behalf of job seekers to vacancies in South Africa. They found that, for the same applicant, attaching a letter increased the probability of receiving a response and the rate of interview requests by about 60 percent.
@@
In summary, I find that prompt-based sentiment extraction produces sentiment scores that are strongly correlated with the probability of obtaining a good job market outcome. In particular, high average sentiment among all referees are predictive of a sizeable higher probability of success. Moreover, my paper is the first to document that disagreement among letter writers can offset the positive effect of high average RL sentiment, which is strictly consistent with employers taking all available information in RLs pretty much at face value. While these findings are not causal, they strongly suggest that the theoretical literature might have over-emphasized the potential pitfalls of RLs.
@@ 2.1 Lexical Approach
A final limitation of the lexical approach is that, if the method is to be used to generate a measure other than a sentiment score, a completely new dictionary must be produced. For example, Schmader et al. (2007) created a set of dictionaries to summarize different semantic aspects of reference letters written for chemists and biochemists. Specifically, they developed sets of terms associated with the following language categories: grindstone traits, ability traits, research terms, teaching and citizenship terms, and standout adjectives. Eberhardt et al. (2023) applied a similar strategy to letters of recommendation written on behalf of job market candidates applying to a top UK economics department.
@@ 2.3.2 Model Choice
All the LLMs mentioned so far belong to the foundational class, meaning that they have not been fine-tuned to follow human instructions or to produce output that is “aligned” with the objectives of humans. Instruct models, like Llama-instruct (Touvron et al., 2023), GPT-instruct (Ouyang et al., 2022), or DeepSeek-V3 (DeepSeek-AI et al., 2025) are “post”-trained using supervised data of question-answer pairs, as well as other techniques like reinforcement learning with human feedback, in which the model produces several alternative outputs that a human user later ranks. [^23] World famous applications like ChatGPT make further additions. For example, all chatbots incorporate a memory layer to the otherwise stateless instruct model.
@@
It is important to reiterate that the “score” that results from these instructions is quite unlike the polarity score from equation (2). The former consists of a sequence of text tokens denoting a floating point number within -1 and 1 that results from the instruct model’s sampling process. The latter is a numerical outcome resulting from summarizing the foundational model’s output vector of probabilities.
An unfortunate implication of the sampling process inherent in text generation by instruct-based LLMs is that it can be difficult to ensure perfect reproducibility of results. Modern APIs allow for setting a seed number, which should in principle ensure reproducibility as long as other parameters (prompt, temperature, model version, etc.) remain unchanged. However, inference providers often update their platforms and this may on occasion affect reproducibility even for same-seed constant-prompt generation. This challenge is altogether absent when using foundational LLMs’ output in a deterministic way.
@@ 3.2 Reference Letter Sentiment
Figure 1 (left) shows the kernel estimate of the distribution of prompt-based polarity scores for the full sample of RLs. Several facts are apparent. Sentiment polarity is overwhelmingly positive; there are only a couple of letters with negative polarity. The distribution of sentiment is roughly symmetric, with mean sentiment of about 0.07.292929Table A.1 in the appendix has descriptive statistics for all the sentiment measures used in this paper.
@@
Letter sentiment tends to reinforce other known signals of higher ability. Candidates from PhD programs in America and from programs in the top 25 have substantially better RLs. The same is true about candidates with top publications, as well as candidates who managed to procure a letter from a top-5 percentile research economist. These statistics are in line with what a general sentiment measure should look like under the general assumption that higher ability candidates get more enthusiastic references.
@@ 4 Predicting Job Market Outcomes
There are multiple potentially appropriate measures of what constitutes “success” in the academic job market. Job market outcomes are heterogenous on multiple fronts. Consider figure 3, depicting different aspects of candidates’ placement by September after the job market season. Outcomes vary by the sector of employment and the type of position. In addition, employers differ on how they are ranked according to their research output, with higher ranked employers being markedly more desirable to the average junior academic. [^30] In order to determine the employer’s ranking, I once again use Tilburg’s ranking for academic institutions. For policy institutions, I use the econphd.net ranking (https://econphd.econwiki.com/rank/rallec.htm), which focuses on publications and has a transparent methodology.
@@
In what follows, I define my baseline “success” outcome as attaining a top-100 academic or top-50 policy job. For some clarity on these cutoffs, consider that economics departments ranked 98–102 in the Tilburg ranking are NHH (Norwegian School of Economics), INSEAD, the California Institute of Technology and the University of Gothenburg. Institutions ranked in the top-50 of the policy ranking include the Federal Reserve Banks and the Board of Governors, as well as the ECB and a few other European national banks, plus the World Bank, the IMF and a couple other major development institutes. In figure 3, the baseline cutoffs are depicted by vertical red lines. Clearly, only a minority of candidates in my sample attained success based on this measure.
@@
Baseline Top-100 Academic Employers & Top-50 Policy Institutions 175 27.1
Alt. 1 Exclude teaching positions (liberal arts colleges) and post-docs from the success group 115 17.8
Alt. 2 Add some highly paid tech jobs (Amazon, Meta, etc.) to the success group 190 29.5
Alt. 3 Use somewhat more forgiving cutoffs: top-150 academic and top-75 policy 243 37.7
Alt. 4 Use somewhat more restrictive cutoffs: top-75 academic and top-35 policy 150 23.3
@@
Notes: The difference in means is $1.33$ p.p. The HC3 standard error is $0.23$.
Figure 4 depicts the relationship between average RL sentiment and baseline success in the academic job market. Clearly, candidates who succeeded to obtain a top job have letters with significantly higher average sentiment.
@@ 4.1 Linear Regressions
$y_{it}=$ $\>\alpha+\beta\,S_{it}+\boldsymbol{X}^{\prime}_{it}\boldsymbol{\gamma}+\epsilon_{it}$ (3)
[…]
RLs are only one input in the recruitment process. Job market outcomes also depend on how the candidates present themselves in interviews, seminar presentations, and one-on-one meetings with faculty and administrators at potential employers, among other variables. Because these other determinants remain unobserved, the regression coefficients cannot be interpreted causally. [^32] Assuming strong candidates get good RLs and also tend to perform well in interviews and seminars, one reasonable expectation would be for regression coefficients to over-estimate the true causal effect that RL sentiment might have on a candidate’s probability of success.
However, in light of the theoretical literature regarding whether referees can be trusted to offer honest assessments and whether hiring committees are properly incentivized to select the best candidates, it is still interesting to determine whether a statistically significant association exists between good letters and good job market outcomes. In this section and the next, I show that not only are higher average sentiment RLs predictive of better outcomes, but also that the degree of disagreement among the letter writers—what I call sentiment dispersion—is negatively associated with success in the market.
@@
Table 7 shows least squares estimates of equation (3) for the baseline outcome and with different sets of control variables. [^33] Table A.2 in the appendix presents the coefficient estimates for the control variables. Regardless of the specifics, RL sentiment has a strong and highly statistically significant relationship with the probability of success. [^34] Throughout the paper I present standard errors, and the corresponding levels of significance, for three different assumptions regarding the joint distribution of the error terms in equation (3). HC3 standard errors assume the $\epsilon_{it}$ are independently distributed but heteroskedastic. Clustered standard errors allow for dependence across errors either within PhD programs (due to norms regarding how to write RLs, for example) or within clusters defined by the market year and the PhD rank group (due to competition among candidates for similar jobs).
Prompt-based sentiment extraction enables a straightforward interpretation of the regression results, even if they cannot be taken to measure a causal effect. For example, an increase in average polarity of 0.1 means that the LLM assigns a 10 pp. higher probability to filling the prompt with one of the positive sentiment tokens relative to the negative tokens. According to the findings in the table, such an increase is predictive of a much higher ($\sim$23 p.p.) expected probability of obtaining a top research job either in academia or at a policy institution.
@@ 4.2 Sentiment Dispersion
An open question is whether disagreement among the many letter writers is reflected in worse job market outcomes for the candidates. In order to study this issue, I focus on a sub-sample of “complete applications,” i.e. candidates for whom I have at least three RLs including the letter from the main research adviser. [^35] Table A.3 in the appendix replicates the analysis in table 8 for this sub-sample. The results are practically the same as with the full sample.
Figure A.1 in the appendix shows that the distributions of sentiment polarity scores for different letter writers are almost identical. [^36] In unreported analysis, I also found that the adviser’s RL sentiment and other writer’s RL sentiment have undistinguishable positive partial correlations with job market outcomes. However, within-candidate variation in RL sentiment is quite high, accounting for over 45% of the total variance of sentiment. [^37] An alternative way to illustrate this point is by looking at correlations across letter writer scores. These correlations range between 0.27 and 0.40. In other words, letter writers tend to be aligned with each other but are far from unanimous.
@@
Notes: NN=553 (complete applications sub-sample). The range is the difference between the $\max$ and the $\min$ sentiment score for each candidate.
Table 9 shows summary statistics for alternative ways to measure sentiment dispersion among letter writers. The sentiment range—the measure that I will use as a baseline—is defined as the difference between the $\max$ and the $\min$ sentiment score among the RLs written on behalf of a candidate. For robustness, I also consider the mean absolute deviation (MAD) and the standard deviation (SD). As is clear from the table, all measures of dispersion are quite high relative to average sentiment. They also vary significantly across candidates.
@@
$y_{it}=$ $\>\alpha+\beta_{1}\,S_{it}+\beta_{2}\,D_{it}+\boldsymbol{X}^{\prime}_{it}\boldsymbol{\gamma}+\epsilon_{it}$ (4)
where $D_{it}$ is one of the measures of sentiment dispersion. Table 10 confirms that, regardless of the outcome definition, sentiment range has a negative relationship to success in the job market. Table A.4 in the appendix shows similar results for alternative measures of sentiment dispersion. In summary, these estimates suggest that hiring committees fully exploit the information contained in RLs.
@@
Four Letters & Avg. RL Length Yes Yes Yes Yes Yes
@@
4.3 Comparison of Prompt-based Sentiment with Other Approaches
Figure 6: Correlation Matrix
Figure 6 presents correlations among the different sentiment measures. While almost all correlations are positive, it is noteworthy that the coefficients are clustered by technique. For example, lexicon-based measures have much stronger associations among each other—correlations close to 0.5—than with any of the LLM-based measures. Similarly, the correlation between the two instruction-based measures is 0.82.
I repeat the regression analysis from section 4.1 for each of the alternative sentiment measures. Figure 7 summarizes my findings. Specifically, the box plots in the figure show how the t-statistic for the coefficient ($\hat{\beta}$) associated with Avg. RL Sentiment in equation (3) varies across specifications. For each of the sentiment measures, I estimated 150 different specifications by varying the outcome variable (5 variants), considering alternative control variable sets (5), using either the full or the restricted ‘complete applications sample’ (2), and applying different assumptions to the error term (3).
Somewhat predictably, neither the lexical approach nor the fine-tuned LLMs are able to predict job market outcomes. The dictionaries involved in the former and the fine-tuning process in the latter target applications that are quite distant from RL sentiment extraction. In fact, in many cases the estimated coefficients obtained using these sentiment measures have the ‘wrong’ sign.
Regardless of these limitations, the sentiment scores that result from applying ModernBert are predictive of job market outcomes. In most of the regressions, the average RL sentiment comes out positive and significant at the 5% level.
Notes: The box plots depict the distributions of t-statistics for the coefficient corresponding to Avg. RL Sentiment in equation (3). Each sentiment measure is used in 150 alternative specifications: 5 outcome variables, 2 samples, 5 control variable groups, and 3 alternative standard errors.
Notes: The box plots depict the distributions of t-statistics for the coefficients corresponding to Avg. RL Sentiment and RL Sentiment Range in equation (4). Each sentiment measure is used in 75 alternative specifications estimated on the ‘full applications’ restricted sample: 5 outcome variables, 5 control variable groups, and 3 alternative standard errors.
Figure 8 repeats the exercise, this time for equation (4). [^39] To save space, I restrict the analysis to the methods that were successful in predicting outcomes without the inclusion of sentiment dispersion. All regressions are estimated on the ‘full application’ restricted sample. Neither ModernBert nor the instruction-based models are able to replicate the findings in table 10 using the Llama 3.1 prompt-based sentiment. Most remarkably, no other sentiment measure is able precisely to capture the role of sentiment dispersion.
4.4 Measurement Error
Our focus of interest is the relationship between RL sentiment and job market outcomes. However, the least squares estimates presented so far are based on a particular approach to measuring sentiment, namely my prompt-based sentiment polarity score. This approach could be subject to bias due to measurement error.
Formally, suppose that the relationship between “true” sentiment $S$ and the prompt-based measure $\hat{S}$ is given by $\hat{S}=S+u$, where $u$ is the measurement error. In general, simply plugging-in the mismeasured variable in equation (3) will lead to least squares estimates that are biased (i.e. they do not consistently estimate the true population linear projection). Without making some assumptions regarding the nature of the error, however, it is not possible to say much else. [^40] Under the classical assumptions ($\mathbf{E}(uS)=\mathbf{E}(u\epsilon)=0$), it is known that the presence of measurement error tends to attenuate the regression coefficients.
Recently, there have been several attempts to analyze the econometric issues that arise in applications that use variables derived from machine learning algorithms or LLMs (e.g. Battaglia et al., 2024, Ludwig et al., 2025). The emerging consensus is that bias introduced by measurement error is indeed the main concern. The remedy suggested in this literature is to de-bias reported estimates. Unfortunately, the bias correction approach requires either a validation sample (a set of data where the variable is observed without error) or detailed knowledge regarding the process generating the error in measurement. This approach is inapplicable for an intrinsically latent variable like RL sentiment ($S$).
A feasible alternative is to take a multiple indicators approach to addressing measurement error. This involves re-estimating the plug-in version of equation (3) by two-stage least squares using alternative measures of sentiment as instruments for $\hat{S}$. Formally, let the instrument be given by $\tilde{S}=S+w$, where $w$ is assumed to be another classical measurement error term. The key assumption necessary for this approach to work is $\mathbf{E}(uw)=0$. In other words, the errors from the two sentiment measures must be uncorrelated. While the assumption is untestable, it is possible to use different instrument candidates with different measurement properties and check for robustness.
Table 11: Two Stage Least Squares Estimates
(1) (2)
Instrument ($\tilde{S}$) ModernBert DeepSeek3 Instruct
First Stage F-stat# 196.0 208.6
Baseline Outcome
[…]
[0.0133]∗∗∗ [0.0111]∗∗∗
[…]
Alternative 1
[…]
Alternative 2
[…]
[0.0140]∗∗∗ [0.0115]∗∗∗
[…]
Alternative 3
[…]
Alternative 4
[…]
[0.0132]∗∗∗ [0.0104]∗∗∗
[…]
Table 11 presents two-stage least squares estimates of equation (3). The first candidate instrument is the prompt-based sentiment polarity measure derived from applying an alternative LLM (ModernBert). On one hand, this sentiment measure applies the same prompt template and the same verbalizer. On the other, it utilizes a bidirectional encoder-based transformer, which uses a different tokenization procedure, a different pre-training technique, and a different training set. [^41] The architectural and methodological differences between ModernBert and Llama 3.1 are very significant. While it is reasonable to assume that much of the raw text that is the input to pre-training is shared, important details regarding the scale, variety, and the kind of pre-processing and cleaning that is applied to the raw text is unique to each project. Moreover, as mentioned above, the language modeling tasks that are the centerpiece of LLM pre-training are very different for encoder and decoders. Finally, the details of the architecture (tokenization approach and vocabulary size, type of positional embeddings, type and placement of normalization layers, type and placement of attention layers, activation functions, etc.) are extremely different. The second instrument is the sentiment score derived from instructing DeepSeek 3. In this case, the model has been post-trained and the prompting strategy is completely different from the one used to obtain my baseline prompt-based sentiment scores. For both instruments, the F-test for the excluded instruments in the first stage regression way exceeds the critical value suggested by Stock and Yogo (2005) to rule out weak instruments.
With almost no exception, all the 2SLS estimates are positive and statistically significant. [^42] This is somewhat remarkable considering the small sample size. Moreover, the 2SLS estimates are sizably larger than the corresponding least squares estimates in table 8. Specifically, the 2SLS estimates are $62\%$ larger than the OLS estimates, on average. Finally, it is noteworthy that the estimates obtained using alternative instruments are numerically quite similar. Taking these results at face value would suggest that the incidence of measurement error is substantial and that the true relationship between RL sentiment and job market outcomes is even stronger than what it might appear based on least squares regressions.